# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ X Closing Remarks
Summarizing, in this paper, I have shown a clear relation between geometry and dynamic features in two-dimensional bidisperse systems of hard disks. This was underpinned by careful analysis of the isoperimetric quotient and mean squared displacement. Surprisingly, the system appears to fall out of equilibrium for a nearly constant value of the area fraction $\phi\approx 0.777$ over a wide range of disk size ratios $R\lesssim 0.83$. I link this feature in the state diagram to the existence of a geometric ground state. Therein the disks are in contact and arranged in such a manner that their Voronoi cells form a floret pentagonal tiling of the plane. This configuration has an area fraction of $\phi_{g}=\sqrt{3}\pi/7\approx 0.777343$, which is in excellent agreement with the observed value of $\phi\approx 0.777$.
I have argued that the existence of such a regular, geometric ground state echoes through in the disordered, finite-temperature system, causing its dynamics to change drastically. This is similar to the way in which the shape index can influence the dynamics in a foam-like tissue models. In terms of the dynamics, the transition picked up in earlier experiments by Lozano et al. [1], has features that are reminiscent of glassy dynamics. However, based on the data presented here, it should not be identified as such. Instead, the viewpoint that I present in this paper, strongly suggests that one should see dense 2D fluids as something else than disordered systems. They may be entangled with ordered, zero-entropy ground states that can enforce geometric frustration even at finite temperature. This perspective could also help explain the observations by Li et al. [2].
Future work in this direction should focus on making the suggested connection mathematically rigorous. I have already briefly discussed extensions to 3D and soft systems in this paper, which provide routes for further exploration of the topic. While presently it appears that the transition obtained in the 2D bidisperse disk systems is not a glass transition, it may also be interesting to see how the concepts introduced here can be modified to help understand structural glass formers.
@@ XI Acknowledgements
The author acknowledges financial support from the Netherlands Organization for Scientific Research (NWO) through Start-Up Grant 740.018.013. I am grateful to Frank Smallenburg for making his code available for use [71], which proved instrumental in the realization of this project, as well as for useful discussions on the analysis. I further acknowledge Edan Lerner for a critical reading of a draft version of the manuscript, which helped me substantially refine my perspective, and Celia Lozano for providing the raw self-intermediate scattering data [1]. Lastly, I would like to thank Luca Giomi, Clemens Bechinger, Roel Dullens, and Alptug Ulugöl for providing useful input on my work.
@@ Appendix A Comparison to the Literature
My study into the dynamics and structure of 2D bidisperse disks builds on decades of research into the structure and dynamics of dense (quasi-)2D suspensions. In this appendix, I will compare my results to those obtained in the literature, where I aim to be as extensive as possible  [^8] It is nearly inevitable that a literature reference has been missed and I apologize should your work unwittingly not receive the same attention as those listed in this section.. Again, referencing the $R$-$\phi$ diagram of Fig. 4, I have put (estimated) data points from the literature on top of my results, labeling these “a” through “q”.
@@
Figure 8: Comparison to the literature. The $R$-$\phi$ state diagram for bidisperse disks in 2D is the same as in Fig. 4a. The dots marked with letters (“a”-“q”; “l” corresponds to the dashed cyan curve) indicate systems for which literature values are known. The large black point “a” represents the experimental system of interest [1] as before. The purple points indicate systems that can be directly mapped to my simulations. Cyan data points represent comparable systems, for which some interpretation of the mapping is required. Brown points follow from theoretical calculations and gray points represent systems that may be compared to my data, but for which the translation is not necessarily obvious. The text in this appendix provides the details on all comparisons and how to interpret these.
@@
Point “a” is the experimental reference [1]. Point “b” is omitted, as I previously discussed this in Fig. 5 and this falls outside of the $\phi$ range shown here. There is also a point “r”, which is discussed in Appendix C, and not shown here to keep the presentation clean. I will now go over these points and discuss how they were obtained and what can be learned from the comparison.
@@
Points “c”-“f”, “j”, “l”, and “m” relate to the rcp value $\phi_{m}$ obtained using the algorithm of Ref. [75]. The first point, labelled “c”, was obtained from the experimental study of Zhao et al. [45], wherein the two-point correlation of free Voronoi volumes in binary disc packings was measured. The size ratio of their teflon disks is $6$ mm to $9$ mm, which results in $R\approx 0.67$. The authors report a maximum value of $\phi=0.838$ for their packing, which is in good agreement with the rcp value that I determined.
@@
From top to bottom, points “d” through “f” correspond to data with $R^{-1}=1.4$ that was obtained from Refs. [110] ($\phi\approx 0.843$), [117] ($\phi\approx 0.8416$), and [118] ($\phi\approx 0.8391$), respectively. The top data point derives from a numerical study into the effect of friction on the jamming properties of disordered particle packings of bidisperse disks [110]. This point was obtained in the frictionless limit, which is appropriate here. The middle data point was extracted from a study of 50:50 bidisperse harmonically repulsive disks with 7:5 diameter ratio [117]. Here, I present the fitted value of the jamming fraction as extracted in Appendix B2 of their work. Note that despite the softness of the potential this extracted jamming point agrees very well with my rcp value. Lastly, the bottom data point specifically examines isostatic, maximally random jammed monodisperse hard-disk packings [118]. This point too matches my $\phi_{m}$ value well, which is expected for a study into jamming.
@@
The point labelled “j”, derives from Ref. [54]. Rieser et al. considered the divergence of Voronoi cell anisotropy vector for both experimental and numerical data on bidisperse disk-like packings [54]. Here, I take the value $\phi\approx 0.8409$ that was obtained using their quenched simulation protocol for $R=0.75$; the $\phi$ for their thermalized data point ($\phi\approx 0.8465$) agreed less well with my rcp value. It may be that thermalization makes the softness of the simulated particles play a more significant role.
@@
The curve labelled “l”, was obtained from Ref. [119], which systematically mapped out the jamming transition of all 2D bidisperse mixtures of frictionless disks in the hard particle limit. In particular, I extracted the data from their Fig. 8 for the largest number of particles simulated and interpolated this to arrive at the dashed cyan curve. Finally, point “m” comes from Ref. [26], wherein a theory of amorphous packings and more generally glassy states of hard spheres is reviewed. Their Table III reports a 2D value of $\phi\approx 0.8745$ that follows from a first-order small cage expansion of the ‘glass close-packing’ density. The agreement is reasonable given the nature of the theoretical approximation employed in Ref. [26].
@@
I conclude that the rcp values obtained using algorithm of Desmond and Weeks [75] agree well with the various pieces of (more modern) literature. Note that I also already commented on the excellent agreement between my numerical result and the experimentally obtained value of $\phi$ for which the SISF relaxation time diverges [1] in Section V. All of this together provides confidence in my choice of setting $\phi_{m}$ as an upper bound to the considered $\phi$ range.
@@
Next, let us cover points “i”, “n”, and “o”, which do not appear to match any feature in the state diagram, before moving onto a discussion of the remaining points that do. Point “i” derives from a numerical study of bidisperse systems ($R^{-1}=1.4$) comprising soft harmonic disks in 2D [120]. The authors report a “thermal jamming transition”, wherein at $\phi\approx 0.74$ an effectively non-ergodic state sets in for which there remain overlaps. The application of this result to the hard-disk packings I have studied is slightly tenuous, however, their $\phi$ value approaches where I place the onset of apparent arrested dynamics. That being said, it much more closely matches the $n_{p}=5.5$ percolation value discussed in Appendix D (not shown here). Thus, there is no obvious conclusion that can be drawn from perceived agreement and any correspondence is more than likely a fortuitous coincidence.
@@
Point “n” was obtained from the combined theorical and experimental work of Ref. [121], wherein an approximate expression for the close-packing density $\phi\approx 0.822$ of a monodisperse sample is proposed. Clearly, the point deviates considerably from the rcp value obtained here, though it agrees with a previous experimental study to within the error [122]. I have nonetheless included this point, as the theory put forward in this source was used to obtain a data point (“p”) that agrees better with one of my measures. Lastly, point “o” follows from the theory of Ref. [26] (see their Table III), and is identified as the point where a thermodynamic glass transition is supposed to happen. Within the limitations of my analysis, I find no evidence for this transition. However, I should stress that I did not pursue analysis of the ideal glass transition; if it is even present for a monodisperse system [96].
@@
This leaves points “g”, “h”, “k”, “p”, and “q”. The first, point “g”, follows from Ref. [96], wherein a binary hard-disk mixture is studied with ratio of disk diameters $R^{-1}=1.4$. The authors report $\phi\approx 0.775$ for the (equilibrium) freezing area fraction. However, this study was performed for a 1:2 stoichiometric ratio of large to small disks. I will argue in Section VIII how this fits my interpretation of $\phi_{a}$.
@@
Next, point “h” follows from Ref. [110], which studied frictional hard-disks with $R^{-1}=1.4$. Here, I fitted the jamming data points with a sigmoidal curve and extrapolated to the infinite friction result to a rlp value of $\phi\approx 0.766$. The reason for taking this fitting approach is that Silbert does not report the 2D jamming value. Moreover, there is a peculiar nonmonotonicity to the coordination-number data for 2D systems provided in Table I of Ref. [110]. Given the level of uncertainty on the data provided, the agreement between my fitted rlp value of $\phi\approx 0.766$ and the arrested-dynamics area fraction $\phi_{a}$ obtained in this study is very good. It is curious that there appears to be a connection between rlp of a fully frictional sample and the apparent arrested dynamics in a frictionless system. My interpretation is that this is another signal of an underlying geometric feature, see Section VIII.
@@
Point “k” was obtained from Ref. [53], the authors of which took a very similar approach to studying 2D systems with arrested dynamics. They considered the evolution of Voronoi cell geometry in simple hard-disk models by simulations and colloid experiments. In particular, they located the ‘glass transition’ using a Voronoi-based (neighborhood) cage anisotropy parameter. Their anisotropy takes the ratio of the long and short axes of the Voronoi cells and has a peak similar to what I find when studying $\max_{\phi}\!\left(n_{\bar{q}}\right)$. Here, I use their point $\phi\approx 0.79$ which was obtained via 2D EDMD of bidisperse disks with $R^{-1}=1.3$ and a 50:50 number ratio. This data point is in good agreement with where I locate $\phi_{a}$ and $\max_{\phi}\!\left(n_{\bar{q}}\right)$, respectively, given the subtle differences in approach and the uncertainty this imposes. However, based on my analysis, a change in cage anisotropy does not fully coincide with the apparent kinetic arrest, though it is possible that a measure based on aspect ratio, as pursued in Ref. [53], leads to a closer connection.
@@
Reference [111] makes use of the theory by Bideau et al. [121] to predict the rlp density of $\phi\approx 0.78$, which is indicated by point “p”, for a monodisperse sample of hard disks. The point “q” follows from the numerical study performed in the same paper, which places the rlp area fraction at $\phi\approx 0.772$. Given the limitations of this study, the agreement with the dashed green line, and thus the minimum in the amount of pentagonal large-particle neighborhoods, is excellent. This is a second piece of evidence — also see point “h” above — in the direction of geometric features underlying the $\phi\approx 0.777$ line in the systems considered in my research. Note that Ref. [111] reports on an average number of contacts per disk being $\approx 3$ at their rlp $\phi$. Using my Voronoi approach, I find an average number of neighbors that is substantially larger, see Fig. 2d. It is, however, interesting to assess how many of these neighbors contribute to ‘immobilizing’ contacts, such as considered in Ref. [111], at any one time.
@@ Appendix B Pseudo-Polygon Fractions
For $R^{-1}>1.35$ (equivalently $R\lesssim 0.74$), I find that there for all $\phi$ I have $f_{6}<f_{5}$ (vertical dotted purple lines in Fig. 9). Note that $R\approx 0.74$ is not value for which the $n_{\bar{q}}=5.5$ curve closes in on itself, which occurs for $R\approx 0.83$. Thus, there is no transition from pentagonal to hexagonal below $R\approx 0.74$, which presumably changes the behavior of the system from being polycrystalline to being truly disordered. However, as is evidenced by the neighborhood composition graphs in Fig. 9, the system takes on a range of compositions as a function of $\phi$ even for relatively low $R$. So one has to be careful in drawing to strong conclusions based on Fig. 9.
@@
Appendix C Pressure
Throughout this paper, I have considered the system in the $NVT$ ensemble (technically at constant area). There is coexistence between a disordered fluid, hexatic phase, and hexagonal crystal at values of $R\approx 1$. Given the system size that I considered, picking up on the hexatic phase is not realistic [123, 124, 125]. However, I should still find the features of a first-order transition between a disordered fluid and ordered hexagonal phase in my data. One such feature is the presence of a van-der-Waals loop in the pressure $p$ as a function of the inverse particle density $v=V/N=\rho^{-1}$. I therefore computed the pressure from the collisions in the system via the momentum exchange during collisions, using the standard virial expression. This was averaged and subsequently used to establish van-der-Waals loops whenever they are present, see Fig. 10a for an example.

### v2

@@ X Closing Remarks
Summarizing, I have argued for the existence of geometric ground states with congruent Voronoi tiles, which induce continuous dynamic transitions in 2D systems of hard disks. The appearance of dynamic arrest then occurs in a manner that is analogous to the one found for model foams and tissues, where geometry and topology are well-known to play a key role.
This insight is based on careful analysis of the isoperimetric quotient, mean squared displacement, and percolation properties. Specifically, the system that I numerically investigated falls out of equilibrium for a nearly constant value of the area fraction $\phi\approx 0.777$ over a wide range of disk size ratios $R\lesssim 0.83$. Together with experimental evidence in both quasi-2D Brownian suspension [1] and 2D granular systems [82, 2, 3], among others, that have dynamical changes for similar values of $\phi$, this points toward a common underlying feature.
I identify this feature to be a geometric ground state, given by a floret pentagonal arrangement of disks on the plane. This configuration has an area fraction of $\phi_{g}=\sqrt{3}\pi/7\approx 0.777343$, which corresponds to the observed transition values. I have tested this argument for a trinary system to validate my approach and found excellent agreement, further suggesting that system composition is not important. In addition, carrying the geometric-ground-state reasoning further, it seems plausible that the experimental observations of a caging transition by Li et al. [4] can be similarly explained using a 2D honeycomb arrangement as geometric ground state. However, the evidence presented here for this correspondence is more tenuous than it is for the glass/random-loose packing transition. In 3D, the reasoning presented here can be carried over to explain the random-loose packing volume fraction for spheres using the square bipyramidal honeycomb arrangement.
Future work in this direction should focus on making the suggested connection between geometry and dynamics more mathematically rigorous. Physically, it is important to probe experimentally the constancy of the transition $\phi$ in both granular and Brownian systems, as well as elucidate which preparations lead to the predicted value of $\phi$ for the former. Thinking in terms of geometric ground states could also help shed light on attempts to explain the crossover between athermal jamming and the thermal glass transition [84, 150], although it should be clear that 3D systems can be quite different from their 2D counterparts.
@@ XI Acknowledgements
The author acknowledges financial support from the Netherlands Organization for Scientific Research (NWO) through Start-Up Grant 740.018.013. I am grateful to Frank Smallenburg for making his code available for use [100], which proved instrumental in the realization of this project, as well as for useful discussions on the analysis. I further acknowledge Clemens Bechinger and Celia Lozano for providing data [1], as well as useful discussion. Lastly, I would like to thank Karen Daniels, Luca Giomi, Wilson Poon, Roel Dullens, Alptug Ulugöl, and Edan Lerner for providing useful input on my work and/or pointing out worthwhile references.
@@ Appendix A Comparison to the Literature
My study into the dynamics and structure of 2D bidisperse disks builds on decades of research into the structure and dynamics of dense (quasi-)2D suspensions. In this appendix, I will compare my results to those obtained in the literature, where I aim to be as extensive as possible  [^13] It is nearly inevitable that a literature reference has been missed; I apologize should your work unwittingly not receive the same attention as those listed in this section.. Again, referencing the $R$-$\phi$ diagram of Fig. 6, I have put (estimated) data points from the literature on top of my results, labeling these “a” through “r”.
@@
Figure 10: Comparison to the literature. The $R$-$\phi$ state diagram for bidisperse disks in 2D is the same as in Fig. 6a. The dots marked with letters (“a”-“r”; “m” corresponds to the dashed cyan curve) indicate systems for which literature values are known. The large black point “a” represents the experimental system of interest [1] and the one marked “b” the one of Refs. [82, 3], as before. The purple points indicate systems that can be directly mapped to my simulations. Cyan data points represent comparable systems, for which some interpretation of the mapping is required. Brown points follow from theoretical calculations and gray points represent systems that may be compared to my data, but for which the translation is not necessarily obvious. The text in this appendix provides the details on all comparisons and how to interpret these.
@@
Point “a” and “b” are the experimental references in thermal [1] and athermal [82, 3], respectively. Point “c” is omitted, as I previously showed this in Fig. 7 and it falls outside of the $\phi$ range shown here. There is also a point “s”, which is covered in Appendix C, and not shown here to keep the presentation clean. I will now go over these points and discuss how they were obtained and what can be learned from the comparison.
@@
Points “d”-“g”, “i”, “m”, and “n” relate to the RCP value $\phi_{m}$ obtained using the algorithm of Ref. [109]. The first point, labelled “d”, was obtained from the experimental study of Zhao et al. [49], wherein the two-point correlation of free Voronoi volumes in binary disc packings was measured. The size ratio of their teflon disks is $6$ mm to $9$ mm, which results in $R\approx 0.67$. The authors report a maximum value of $\phi=0.838$ for their packing, which is in good agreement with the RCP value that I determined.
@@
From top to bottom, points “e” through “g” correspond to data with $R^{-1}=1.4$ that was obtained from Refs. [135] ($\phi\approx 0.843$), [159] ($\phi\approx 0.8416$), and [160] ($\phi\approx 0.8391$), respectively. The top data point derives from a numerical study into the effect of friction on the jamming properties of disordered particle packings of bidisperse disks [135]. This point was obtained in the frictionless limit, which is appropriate here. The middle data point was extracted from a study of 1:1 bidisperse harmonically repulsive disks with 7:5 diameter ratio [159]. Here, I present the fitted value of the jamming fraction as extracted in Appendix B2 of their work. Note that despite the softness of the potential this extracted jamming point agrees very well with the RCP value reported here. Lastly, the bottom data point specifically examines isostatic, maximally random jammed monodisperse hard-disk packings [160]. This point too matches my $\phi_{m}$ value well, which is expected for a study into jamming.
@@
The point labelled “i”, derives from Ref. [58]. Rieser et al. considered the divergence of Voronoi cell anisotropy vector for both experimental and numerical data on bidisperse disk-like packings [58]. Here, I take the value $\phi\approx 0.8409$ that was obtained using their quenched simulation protocol for $R=0.75$; the $\phi$ for their thermalized data point ($\phi\approx 0.8465$) agreed less well with the RCP value in Fig. 10. It may be that thermalization makes the softness of the simulated particles play a more significant role.
@@
The curve labelled “m”, was obtained from Ref. [161], which systematically mapped out the jamming transition of all 2D bidisperse mixtures of frictionless disks in the hard particle limit. In particular, I extracted the data from their Fig. 8 for the largest number of particles simulated and interpolated this to arrive at the dashed cyan curve. Finally, point “n” comes from Ref. [30], wherein a theory of amorphous packings and more generally glassy states of hard spheres is reviewed. Their Table III reports a 2D value of $\phi\approx 0.8745$ that follows from a first-order small cage expansion of the ‘glass close-packing’ density. The agreement is reasonable given the nature of the theoretical approximation employed in Ref. [30].
@@
I conclude that the RCP values obtained using algorithm of Desmond and Weeks [109] agree well with the various pieces of (more modern) literature. Note that I also already commented on the excellent agreement between my numerical result and the experimentally obtained value of $\phi$ for which the SISF relaxation time diverges [1] in Section V. All of this together provides confidence in my choice of setting $\phi_{m}$ as an upper bound to the $\phi$ range.
@@
Next, let us cover points “j”, “o”, and “p”, which do not appear to match any feature in the state diagram, before moving onto a discussion of the remaining points that do. Point “j” derives from a numerical study of bidisperse systems ($R^{-1}=1.4$) comprising soft harmonic disks in 2D [162]. The authors report a “thermal jamming transition”, wherein at $\phi\approx 0.74$ an effectively non-ergodic state sets in for which there remain overlaps. The application of this result to the hard-disk packings I have studied is slightly tenuous, however, their $\phi$ value approaches where I place the onset of apparent arrested dynamics. That being said, it much more closely matches the $n_{p}=5.5$ percolation value discussed in Appendix D (not shown here). Thus, there is no obvious conclusion that can be drawn from perceived agreement and any correspondence is more than likely a fortuitous coincidence.
@@
Point “o” was obtained from the combined theorical and experimental work of Ref. [137], wherein an approximate expression for the close-packing density $\phi\approx 0.822$ of a monodisperse sample is proposed. Clearly, the point deviates considerably from the RCP value obtained here, though it agrees with a previous experimental study to within the error [146]. I have nonetheless included this point, as the theory put forward in this source was used to obtain a data point (“q”) that agrees better with one of my measures. Lastly, point “p” follows from the theory of Ref. [30] (see their Table III), and is identified as the point where a thermodynamic glass transition is supposed to happen. Within the limitations of my analysis, I find no evidence for this transition. However, I should stress that I did not pursue analysis of the ideal glass transition; if it is even present for a monodisperse system [133].
@@
This leaves points “h”, “i”, “l”, “q”, and “r”. The first, point “h”, follows from Ref. [133], wherein a binary hard-disk mixture is studied with ratio of disk diameters $R^{-1}=1.4$. The authors report $\phi\approx 0.775$ for the (equilibrium) freezing area fraction. However, this study was performed for a 1:2 stoichiometric ratio of large to small disks. I will argue in Section VIII how this fits my interpretation of $\phi_{a}$.
@@
Next, point “i” follows from Ref. [135], which studied frictional hard-disks with $R^{-1}=1.4$. Here, I fitted the jamming data points with a sigmoidal curve and extrapolated to the infinite friction result to a RLP value of $\phi\approx 0.766$. The reason for taking this fitting approach is that Silbert does not report the 2D jamming value. Moreover, there is a peculiar nonmonotonicity to the coordination-number data for 2D systems provided in Table I of Ref. [135]. Given the level of uncertainty on the data provided, the agreement between my fitted RLP value of $\phi\approx 0.766$ and the arrested-dynamics area fraction $\phi_{a}$ obtained in this study is very good. It is curious that there appears to be a connection between RLP of a fully frictional sample and the apparent arrested dynamics in a frictionless system. My interpretation is that this is another signal of an underlying geometric feature, see Section VIII.
@@
Point “l” was obtained from Ref. [57], the authors of which took a very similar approach to studying 2D systems with arrested dynamics. They considered the evolution of Voronoi cell geometry in simple hard-disk models by simulations and colloid experiments. In particular, they located the glass transition using a Voronoi-based (neighborhood) cage anisotropy parameter. Their anisotropy takes the ratio of the long and short axes of the Voronoi cells and has a peak similar to what I find when studying $\max_{\phi}\!\left(n_{\bar{q}}\right)$. Here, I use their point $\phi\approx 0.79$ which was obtained via 2D EDMD of bidisperse disks with $R^{-1}=1.3$ and a 1:1 number ratio. This data point is in good agreement with where I locate $\phi_{a}$ and $\max_{\phi}\!\left(n_{\bar{q}}\right)$, respectively, given the subtle differences in approach and the uncertainty this imposes. However, based on my analysis, a change in cage anisotropy does not fully coincide with the apparent kinetic arrest, though it is possible that a measure based on aspect ratio, as pursued in Ref. [57], leads to a closer connection.
@@
Reference [136] makes use of the theory by Bideau et al. [137] to predict the RLP density of $\phi\approx 0.78$, which is indicated by point “q”, for a monodisperse sample of hard disks. The point “r” follows from the numerical study performed in the same paper, which places the RLP area fraction at $\phi\approx 0.772$. Given the limitations of this study, the agreement with the dashed green line, and thus the minimum in the amount of pentagonal large-particle neighborhoods, is excellent. This is a second piece of evidence — also see point “i” above — in the direction of geometric features underlying the $\phi\approx 0.777$ line in the systems considered in my research. Note that Ref. [136] reports on an average number of contacts per disk being $\approx 3$ at their RLP $\phi$. Using my Voronoi approach, I find an average number of neighbors that is substantially larger, see Fig. 4d. It is, however, interesting to assess how many of these neighbors contribute to ‘immobilizing’ contacts, such as considered in Ref. [136], at any one time.
Lastly, a recent work on the ideals glass transition in 2D disks for a polydisperse systems reports $\phi\approx 0.78$ [138], which is again close to the FPT value that I obtained. The polydisperse nature of the system provides further (circumstantial) evidence for an underlying geometric feature to the transition.
@@ Appendix B Pseudo-Polygon Fractions
For $R^{-1}>1.35$ (equivalently $R\lesssim 0.74$), I find that for all $\phi$ the inequality $f_{6}<f_{5}$ holds (vertical dotted purple lines in Fig. 11). Note that $R\approx 0.74$ is not value for which the $n_{\bar{q}}=5.5$ curve closes in on itself, which occurs for $R\approx 0.83$. Thus, there is no transition from pentagonal to hexagonal below $R\approx 0.74$, which presumably changes the behavior of the system from being polycrystalline to being truly disordered. However, as is evidenced by the neighborhood composition graphs in Fig. 11, the system takes on a range of compositions as a function of $\phi$ even for relatively low $R$. So one has to be careful in drawing to strong conclusions based on Fig. 11.
@@
Appendix C Crystallinity and Pressure
In this section, the pressure and bond-orientational order parameter are considered. Figure 12a compares the $n_{\bar{q}}=5.5$ curve and the behavior of the traditional Steinhardt order parameter $\psi_{6}$. Smoothing and interpolation had to be applied to achieve the surface plot, see the ESI [64] for details. The literature provides reasonable average values for $\psi_{6}$ [163], which the data in Fig. 12a adheres to. However, to the best of my knowledge, clear dividing values for fluid and (locally) crystalline configurations, as are available for 3D, are lacking. Here, I assume that $\psi_{6}\approx 0.54$ is a good divide for when crystalization occurs, based on the $\phi$ values for which a monodisperse system is known to crystallize. This gives agreeable correspondence with the $n_{\bar{q}}=5.5$ curve.
The is coexistence between a disordered fluid, hexatic phase, and hexagonal crystal at values of $R\approx 1$ has attracted considerable historic interest [126, 127, 121, 122, 128]. Given the system size that I considered, picking up on the hexatic phase is not realistic [121, 122]. However, there should be features of a first-order transition between a disordered fluid and ordered hexagonal phase in my data.